OpenStudy (anonymous):

A region is 1.5 cm in length and the electrons enter with a speed of 1x10^5 m/s and leaves with a speed of 2.5x10^6 m/s What is their acceleration over this 1.5 cm length?

OpenStudy (anonymous):

Using the equation of motion displacement, initial velocity and final velocity are given, you can now find the acceleration

OpenStudy (anonymous):

The equation of motion? Doesn't that need time?

OpenStudy (anonymous):

isn't there any equation of motion in which there is no time?

OpenStudy (***[isuru]***):

Hi, I think u guys have missed the equation v^2 = u^2 + 2aS where v = final velocity u = initial velocity a = acceleration S = Distance/Displacement In ur case u = 1x10^5 m/s v = 2.5x10^6 m/s S = 1.5 cm = 0.015 m a = it's the question Now all we have to do is just substituting the values in the equation v^2 = u^2 +2aS (2.5x10^6)^2 = (1x10^5)^2 + (2*0.015*a) (2*0.015*a) = 10^10[ ( 6.25*10^2 - 1] (2*0.015*a) = 624*10^10 a = 624*10^10/(2*0.015) = 312*10^10 / (1.5*10^-2) = 208 * 10^12 m/s^2 That's the answer for ur question acceleration = 208 * 10^12 m/s^2 Hope this will help u !!!

OpenStudy (anonymous):

I tried that answer but it said it was wrong? Is there another possible way to do it?

OpenStudy (anonymous):

Oh wait never mind thanks for the help :) how do you determine how long it's been in the accelerating region?

OpenStudy (***[isuru]***):

Did u say "it said it was wrong" ? Then it seems that u know something or someone who knows the correct answer. It would be really helpful if u could reply the answer 'cause then I could find out where I have messed up If u want to find the time period of acceleration S = ( u + v ) / t (where the signs stand as same in the above one) 1.5 * 10^-2 = [(1x10^5) + (2.5x10^6)]/2 * t t = 1.5/13*10^7 s and u know u could substitute it in S = ut + 1/2at^2 and find "a" but for my amazement it still gives me the same answer

OpenStudy (anonymous):

Ya it's an computer that needs the correct answer but it was just a matter of plugging it in correctly, your answer is right :)